# Combinatorial Selection with Costly Information

### v1

@@ 1 Introduction
  • Pandora’s Box with Partial Inspection (Section 5). This setting is an extension of the classical PB where in addition to opening the box, the algorithm can choose to “peek” into the box at a cheaper cost and learn its value; boxes must still be opened before being selected. We consider the minimization version of this problem and present a $\sqrt{2}$-approximation using the notion of local approximation. The minimization version of this problem has not been studied previously.
  • Weighing Scale Problem (Section 6). This is the problem described in the introduction, and is a new problem introduced in this paper. We obtain a logarithmic approximation in key parameters of the problem using the notion of pointwise approximation.
  • Pandora’s Box with Optional Inspection (Section 7). This setting is an extension of the classical PB where the algorithm is allowed to select boxes without opening them. We consider the maximization version of this problem. We show that no policy can achieve better than a $0.5$-local approximation. On the other hand, semilocal approximation allows us to achieve a $0.582$-approximation. The maximization version of the problem is extensively studied but previous results only apply to the single selection setting. Our result, in contrast, extends to matroid feasibility constraints.
Our cost amortization framework is presented in Section 3 and local approximation in Section 4.
1.1 Related work
Bandit superprocesses (a.k.a. BSPs) were first defined by Nash (1973). Gittins (1979) first showed that optimal policies for the MAB are indexable and introduced the Gittins index (see, also, Nash (1980); Weber (1992)). Weitzman (1979) independently proved the optimality of index policies for the Pandora’s Box special case. Further progress on combinatorial selection over Markov chains was made by Dumitriu et al. (2003); Singla (2017) and Gupta et al. (2019), with the last paper showing the approximate optimality of indexing-based frugal algorithms for this setting. For general BSPs, Whittle (1980) provided an alternate proof of the Gittins index theorem and introduced optimality curves and the Whittle integral, extending the optimality of indexed policies to BSPs with dominant local policies (known in the literature as “Whittle’s condition”). Glazebrook (1982) showed that BSPs are not indexable in general. Several decades later Brown and Smith (2013) proved that the Whittle integral provides an upper bound on the optimal value for a general maximization BSP; their work applies to the single selection problem in the infinite horizon discounted setting and is based on dynamic programming. Our amortization-based bound is identical to the Whittle integral but is algorithmic and constructive; applies to the finite horizon setting as well as to combinatorial selection; and, as a side product, provides surrogate costs that prove useful in designing approximations.
@@ 1.1 Related work
There is little work on BSPs that do not satisy Whittle’s condition, outside of the variants of Pandora’s Box. The only such work we are aware of is by Ke and Villas-Boas (2019), who look at a problem with two specific symmetric MDPs and an arbitrary outside option, and develop an adaptive (but complicated) exact algorithm for this problem. We refer the reader to the survey by Hadfield-Menell and Russell (2015) for further discussion of BSPs.
@@
The idea of augmenting the inspection process of a Pandora’s box in order to explore more interesting decision-making settings has been a very active line of research over the last years. Most literature on the optional inspection variant of the problem is for single-item selection, as opposed to combinatorial variants. Study of it was initiated by Guha et al. (2008), who give a $4/5$-approximation for the maximization setting; and Doval (2018), who characterized the solution to the single-box problem and proved certain conditions under which the Gittins policy remains optimal for single-item selection in the maximization setting (though the results naturally extend to the minimization setting). Beyond this results are separated by whether they are for the minimization or maximization setting. For the minimization setting, Scully and Doval (2024) proved a composition theorem for a special case of local approximation (Definition 7) and used it to construct a committing policy with a $4/3$-approximation guarantee. This result extends as-is to the combinatorial setting via Singla (2017)’s frugal algorithms framework.
In the maximization setting, Beyhaghi and Kleinberg (2019) and Guha et al. (2008) give approximation guarantees for committing policies. Furthermore, Fu et al. (2022) and Beyhaghi and Cai (2022) introduced polynomial time approximation schemes that for any $\varepsilon>0$ compute a policy that is at least a $(1-\varepsilon)$-approximation. However, all of these results are for the single-item selection setting and do not extend to combinatorial settings such as matroid selection. Prior to this work, no policy was known to obtain better than a straightforward $1/2$-approximation for matroid selection PBOI. We show in Section 7 that while local approximation is inadequate for beating the $1/2$, our novel approach of semilocal approximation provides a $0.582$ approximation.
@@
The partial inspection variant of Pandora’s Box has primarily been studied in the context of maximization. Aouad et al. (2020) provide a $(1-1/e)$-approximation via a committing policy, and show that, in fact, any committing algorithm or its negation (flipping which box should be partially opened versus fully opened) admits a $(1/2)$-approximation to the optimal utility. We note that this already highlights a significant difference between the minimization and maximization settings for this variant. Whereas for maximization one can essentially flip a coin to decide which of the two actions to commit to, obtaining the optimal’s utility for the committed action and non-negative utility for the action that was not selected, the same approach cannot be applied for minimization as the cost suffered by a bad flip could result to arbitrarily bad approximations. Beyhaghi (2019) introduces a more general inspection model, where the searcher has $k$ different methods for inspecting each box, and can select at most one of them. They provide a $(1-1/e)$-approximation that applies to $k$-element selection, but is computationally inefficient when $k$ is large. To our knowledge, partial inspection has not been studied in the context of minimization.
@@
Other extensions of Pandora’s Box include settings with combinatorial rewards (Olszewski and Weber, 2015); combinatorial costs (Berger et al., 2023); correlated values (Chawla et al., 2019, 2021; Gergatsouli and Tzamos, 2024); or constraints on the order of inspection (Esfandiari et al., 2019; Boodaghians et al., 2020; Bowers and Waggoner, 2024). We refer the reader to the survey by Beyhaghi and Cai (2024) for further discussion of Pandora’s Box.
@@
The weighing scale problem has not been studied previously, although Hoefer et al. (2024) study a similar setting, where each alternative can be weighed against the median of its distribution (conditioned on any past weighings). They consider a model where the algorithm is provided a budget on the number of weighings and wants to find the best alternative subject to the budget; and design a constant factor approximation in this setting.
2 Preliminaries
@@
We begin by defining a single component of our algorithmic problem that involves learning information about a random variable through a series of costly steps. We represent this process as a Markov Decision Process (henceforth, MDP), instantiated over this random variable. An MDP is described by specifying its state space, action set and how states transition upon taking actions. In the case of Costly Information MDPs, states represent the information the algorithm has gained about the corresponding random variable. Accordingly, we associate each state with the conditional value distribution it represents. Formally:
Definition 1.
@@
Furthermore, $V(\cdot)$ is a function mapping states to distributions over values. $V(\sigma)$ is the (unconditional) distribution of $X$ and $V(s)$ is the posterior distribution of $X$ conditioned on being in state $s\in S$. As such, $V$ satisfies the rules of conditional probability: for all states $s\in S$ and all actions $a\in A(s)$, we have $V(s)=\sum_{s^{\prime}}\mathcal{D}(s,a,s^{\prime})V(s^{\prime})$. We also write $v(s):=\operatorname{E}\left[V(s)\right]$. Finally, $T\subseteq S$ is the set of terminal states. Terminal states have only one action available, called the “accept” action. This accept action comes at no cost; results in a value of $v(s)$ at terminal state $s\in T$; and terminates the MDP.
We call the process $\mathcal{M}_{X}$ a Costly Information Markov chain if there is only one action (accept or advance) available at every state $s\in S$. When clear from the context, we will drop the subscript $X$. For simplicity, we will often refer to these components simply as “MDPs” and “Markov chains”.
At the expense of blowing up the size of the state space and action set, we will assume that the sets of actions $A(s)$ and $A(s^{\prime})$ are disjoint for $s\neq s^{\prime}$, and furthermore that each state in the MDP $\mathcal{M}_{X}$ is reached via a unique sequence of actions starting from $\sigma$; in other words, the state “memorizes” the sequence of actions taken to reach it and thus our MDPs have a tree-like structure. We also assume that each MDP $\mathcal{M}_{X}$ has a finite horizon: there is a constant $H$ such that any state reached via a sequence of $H$ steps is a terminal state. Finally, for simpler exposition, we will assume that the state spaces, action sets and the supports of the random variables $X$ are finite. We note that our entire framework seamlessly extends to continuous settings.
We are now ready to define selection problems with costly information:
Definition 2.
@@
The Costly Information Combinatorial Selection problem (henceforth, CICS) is defined over a ground set of $n$ random variables, $X_{1},X_{2},\cdots,X_{n}$; a feasibility constraint $\mathcal{F}\subseteq 2^{[n]}$; and a costly information MDP $\mathcal{M}_{i}:=\mathcal{M}_{X_{i}}=(S_{i},\sigma_{i},A_{i},c_{i},\mathcal{D}_{i},V_{i},T_{i})$ for each variable $X_{i}$. The constraint $\mathcal{F}$ corresponds to an upwards closed set in the minimization version (henceforth, min-CICS), and to a downwards closed set in the maximization version (henceforth, max-CICS).
@@
An algorithm (a.k.a. adaptive policy) for CICS proceeds as follows. Let $S\subseteq[n]$ denote the set of indices of all terminated MDPs. Let $s_{i}$ denote the current state of the MDP $\mathcal{M}_{i}$ at any point of time during the process. Initially $S=\emptyset$ and $s_{i}=\sigma_{i}$ for all $i\in[n]$. The algorithm chooses at every step an index $i\in[n]\setminus S$ corresponding to a non-terminated MDP $\mathcal{M}_{i}$ and an action $a_{i}\in A_{i}(s_{i})$. It then follows the action at a cost of $c_{i}(a_{i})$. If $a_{i}$ is the accept action (i.e. $s_{i}\in T_{i}$ is a terminal state), it adds $i$ to $S$ and collects the value $v_{i}(s_{i})$. Otherwise, it updates the state of $\mathcal{M}_{i}$ to a new state drawn from the distribution $\mathcal{D}(s_{i},a_{i},\cdot)$, while the states of all other MDPs $\mathcal{M}_{i^{\prime}}$ with $i^{\prime}\neq i$ remain unchanged. Observe that the algorithm can make both of these choices – the index of the MDP to move in and the action to take in that MDP – adaptively depending on the evolution of all $n$ MDPs in previous steps.
@@
Feasibility Constraints.
We focus primarily on matroid feasibility constraints: each MDP corresponds to an element of some known matroid $\mathbb{M}=([n],\mathcal{I})$. For min-CICS, $\mathcal{F}$ is the collection of all sets that contain a basis of $\mathbb{M}$; for max-CICS, $\mathcal{F}$ contains all the independent sets of $\mathbb{M}$. Observe that single-element selection, where the algorithm’s goal is to accept one MDP (e.g. in the classic version of Pandora’s Box or the weighing scale problem described in the introduction), is a special case of matroid selection.
Committing Algorithms.
A committing policy $\pi$ for an MDP $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ maps every state $s\in S$ to a distribution $\pi(s)$ over actions in $A(s)$. An MDP $\mathcal{M}$ coupled with a committing policy $\pi$ defines a Costly Information Markov chain, denoted $\mathcal{M}^{\pi}$, where the single action available in every state is given by the distribution over actions specified by $\pi$. Observe that the states in $\mathcal{M}^{\pi}$ are a subset of the states in $\mathcal{M}$ and terminal states in $\mathcal{M}$ continue to be terminal states in $\mathcal{M}^{\pi}$. We use $\mathcal{C}\left(\mathcal{M}\right)$ to denote the set of all committing policies $\pi$ for an MDP $\mathcal{M}$.
A commitment for a CICS is defined as a tuple $\Pi=(\pi_{1},\cdots,\pi_{n})$ of committing policies $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$, one for each of the $n$ MDPs comprising the CICS. A committing algorithm under $\Pi$ chooses a feasible index $i$ at every step, samples an action from the distribution $\pi_{i}(s_{i})$ and plays that action. Observe that although the action taken by the algorithm is pre-chosen in every state, the algorithm can choose the index of the MDP to move in at every step adaptively based on the evolution of all $n$ MDPs in previous steps. Committing algorithms are therefore adaptive algorithms, although they are a simpler and smaller class relative to the class of all algorithms.
Henceforth, we will focus primarily on matroid-min-CICS. We detail the changes needed to the framework for the maximization setting in Appendix E. Following the work of Singla (2017), our results will apply to any combinatorial setting for which the underlying constraint admits an efficient frugal approximation algorithm. We describe this extension in Appendix F.

### v2

@@ 1.1 Our technical contributions and their relationship to prior work
  • PB with Partial Inspection. In addition to opening the box, the algorithm can choose to “peek” into it at a cheaper cost and learn its value; boxes must still be opened before being selected. We prove that the commitment gap is at most $\sqrt{2}$.
  • Additive PB. The value of the box is given by the sum of independent random variables; each component must be separately inspected at a cost and the box can be selected only after all components have been inspected. This models the shortest paths problem over disjoint paths described earlier. We prove that the commitment gap is at most $2$.
  • Weighing Scale Problem. The box cannot be opened, but can be compared at a cost against a threshold $t$ (chosen by the algorithm); the comparison reveals whether $X_{i}\geq t$ or not. Any sequence and number of comparisons are allowed. We obtain a logarithmic bound (in key parameters of the problem) on the commitment gap.
All these results apply to any set of matroid feasibility constraints, and can be seamlessly extended to any set of constraints that admit a frugal algorithm, as defined by Singla (2017). Furthermore, they immediately extend to settings that consider combinations of these variants: an instance consisting of some (normal) boxes, some partial inspection boxes and some additive boxes under matroid feasibility constraints still has a commitment gap of at most $2$.
Furthermore, we prove that the commitment gap of PB with Optional Inspection under matroid feasibility constraints in the maximization setting is at least $0.58$. This problem has been extensively studied in the literature; for the single-selection setting, it is known that the problem admits a PTAS (Fu et al., 2022; Beyhaghi and Cai, 2022). For matroid constraints, Beyhaghi and Kleinberg (2019) show that the commitment gap is at least $0.63$; however, their approach cannot be used to determine the committing policy that achieves it, as it requires an enumeration of the exponentially many committing policies. In contrast, the committing policy achieving our bound can be efficiently computed and leads to the first approximation result that is strictly better than the (trivial) bound of $0.5$.
1.2 Further related work
@@
We have already discussed the relationship of CICS to bandit superprocesses and Markovian multi-armed bandit. There is little work on BSPs that do not satisy Whittle’s “unambiguous local optimality” condition, outside of the variants of Pandora’s Box. The only such work we are aware of is by Ke and Villas-Boas (2019), who look at a problem with two specific symmetric MDPs and an arbitrary outside option, and develop an adaptive (but complicated) exact algorithm for this problem. We refer the reader to the survey by Hadfield-Menell and Russell (2015) for further discussion of BSPs.
Combinatorial variants of Pandora’s Box were first studied by Singla (2017) who showed that for feasibility constraints admitting greedy-style or “frugal” approximation algorithms, an extension of Weitzman’s indexing algorithm provides the same approximation factor. This was further extended to CICS over Markov chains by Gupta et al. (2019).
@@
The idea of augmenting the inspection process of a Pandora’s box in order to explore more interesting decision-making settings has been a very active line of research over the last years. Arguably, the most studied variant is optional inspection. Most literature focuses on the single-item selection, as opposed to combinatorial variants. Study of it was initiated by Guha et al. (2008), who give a $4/5$-approximation for the maximization setting; and Doval (2018), who characterized the solution to the single-box problem and proved certain conditions under which the Gittins policy remains optimal for single-item selection in the maximization setting (though the results naturally extend to the minimization setting). Beyond this results are separated by whether they are for the minimization or maximization setting. For the minimization setting, Scully and Doval (2024) proved a composition theorem for a special case of local approximation (Definition 8) and used it to construct a committing policy with a $4/3$-approximation guarantee. In fact, their result extends beyond the single-item selection setting to the combinatorial Pandora’s box setting studied by Singla (2017). In the maximization setting, Beyhaghi and Kleinberg (2019) and Guha et al. (2008) give approximation guarantees for committing policies. Furthermore, Fu et al. (2022) and Beyhaghi and Cai (2022) introduced polynomial time approximation schemes that for any $\varepsilon>0$ compute a policy that is at least a $(1-\varepsilon)$-approximation. However, all of these results are for the single-item selection setting. For matroid feasibility constraints, Beyhaghi and Kleinberg (2019) observe that the commitment gap is at least $(1-1/e)$ and give an efficient policy that achieves a $0.5$ approximation.
@@
The partial inspection variant of Pandora’s Box has also primarily been studied in the context of maximization. Aouad et al. (2020) provide a $(1-1/e)$-approximation via a committing policy, and show that, in fact, any committing algorithm or its negation (flipping which box should be partially opened versus fully opened) admits a $(1/2)$-approximation to the optimal utility. We note that this already highlights a significant difference between the minimization and maximization settings for this variant. Whereas for maximization one can essentially flip a coin to decide which of the two actions to commit to, obtaining the optimal’s utility for the committed action and non-negative utility for the action that was not selected, the same approach cannot be applied for minimization as the cost suffered by a bad flip could result to arbitrarily bad approximations. Beyhaghi (2019) introduces a more general inspection model, where the searcher has $k$ different methods for inspecting each box, and can select at most one of them. They provide a $(1-1/e)$-approximation that applies to $k$-element selection, but is computationally inefficient when $k$ is large. To our knowledge, partial inspection has not been studied in the context of minimization.
The additive PB model has been studied previously (Bowers and Waggoner, 2024a, b) in the context of maximization for $k=2$ random variables per alternative with matching constraints. In this case, each alternative admits only two commitments: opening one of the two random variables followed by the other. The authors show via a global argument that picking one of the two commitments randomly independently for each alternative yields a bound of $2$ on the commitment gap (a further factor of $2$ is lost due to the matching constraint). For boxes with $k>2$ components each, applying the same approach implies a commitment gap of $O(k!)$. We leave as an open question whether a better bound can be obtained using our techniques. The minimization variant of this problem has not been explored in prior research.
It is noteworthy that a common approach in the literature to bounding the commitment gap in maximization settings involves reducing the problem to finding the optimal committing policy for a stochastic submodular maximization instance and applying the bound of Asadpour and Nazerzadeh (2016) on the adaptivity gap for such problems (Beyhaghi and Kleinberg, 2019; Aouad et al., 2020; Beyhaghi, 2019). This method is inherently global, as it requires efficient optimization over the entire instance once mapped to a stochastic submodular maximization framework. Moreover, the optimization scales with the number of committing policies and does not extend beyond the single-selection case.
@@
Other extensions of Pandora’s Box include settings with combinatorial rewards (Olszewski and Weber, 2015); combinatorial costs (Berger et al., 2023); correlated values (Chawla et al., 2019, 2021; Gergatsouli and Tzamos, 2024); or constraints on the order of inspection (Esfandiari et al., 2019; Boodaghians et al., 2020; Bowers and Waggoner, 2024a). We refer the reader to the survey by Beyhaghi and Cai (2024) for further discussion of Pandora’s Box.
@@
The weighing scale problem has not been studied previously, although similar settings where the algorithm is provided with a budget on the total number of weighings it can perform have been considered. In particular, Hoefer et al. (2024) study a setting, where each alternative can be weighed against the median (or any other fixed quantile) of its distribution (conditioned on any past weighings). Hoefer and Schewior (2023) also consider the setting where the alternatives are identically distributed and the decision-maker is allowed to weigh them against any threshold of their choosing (much like the weighing scale problem). In both of these works, using the weighing scale is free; the algorithm is provided with a budget on the number of weighings it can perform; the objective is to find the best alternative subject to the budget; and constant factor approximations are obtained.
Simultaneous Work.
In simultaneous and independent work, Bowers et al. (2025) also study CICS in the maximization setting. Rather than approaching the problem through the study of the commitment gap, they consider an ex-ante relaxation of the optimal policy that allows them to efficiently obtain a non-committing approximation policy in settings where the underlying combinatorial constraint admits a prophet inequality; in particular, they obtain a $0.5$-approximation under matroid constraints. A key difference between our frameworks is that (i) our results and local approximation also apply to the minimization setting, and (ii) the bounds we obtain on the commitment gap also depend on the structure of the underlying MDPs, allowing for improved approximations in some cases (such as PB with Optional Inspection).
1.3 Outline of the rest of the paper
In Section 2, we provide formal definitions for CICS and the commitment gap. Our amortization framework and bound on the optimal policy for the minimization setting are presented in Section 3; corresponding results for the maximization setting are established in Appendix A. The local approximation framework and its composition to bounds for the commitment gap are presented in Section 4. Subsequent sections are dedicated to bounding the commitment gap for different instantiations of CICS: PB with Partial Inspection (Section 5), Additive PB (Section 6), Weighing Scale Problem (Section 7) and Optional Inspection PB (Section 8).
2 Definitions
@@
We begin by defining costly information acquisition for a random variable as a Markov decision process (MDP). In a Costly Information MDP, states represent the information the algorithm has gained about the corresponding random variable. Accordingly, we associate each state with the conditional value distribution it represents. Formally:
Definition 1 (Costly Information MDP).
@@
$V(\cdot)$ is a function mapping states to distributions over values. $V(\sigma)$ is the (unconditional) distribution of $X$ and $V(s)$ is the posterior distribution of $X$ conditioned on being in state $s\in S$. As such, $V$ satisfies the rules of conditional probability: for all states $s\in S$ and all actions $a\in A(s)$, we have $V(s)=\sum_{s^{\prime}}\mathcal{D}(s,a,s^{\prime})V(s^{\prime})$. We also write $v(s):=\operatorname{E}\left[V(s)\right]$. Finally, $T\subseteq S$ is the set of terminal states. Terminal states have only one action available, called the “accept” action. This accept action comes at no cost; results in a value of $v(s)$ at terminal state $s\in T$; and terminates the MDP. In the special case where there is only one action (accept or advance) available at every state $s\in S$, we call $\mathcal{M}_{X}$ a Costly Information Markov chain. When clear from the context, we will drop the subscript $X$.
Assumptions.
Note that each costly information MDP $\mathcal{M}_{X}$ is inherently acyclic (i.e. the underlying graph is a DAG), as each costly action further refines the algorithm’s knowledge of the underlying random variable. For simplicity of exposition and without loss of generality, we assume that the state spaces, action sets, and the support of the random variable $X$ are finite, although our framework and results extend seamlessly to continuous settings. We will also assume a bounded horizon. In particular, there exists a constant $H$ such that any state reached after a sequence of $H$ actions is terminal.
Definition 2 (Costly Information Combinatorial Selection).
@@
The Costly Information Combinatorial Selection problem (henceforth, CICS) is defined over a ground set of $n$ nonnegative random variables, $X_{1},X_{2},\cdots,X_{n}$; a costly information MDP $\mathcal{M}_{i}:=\mathcal{M}_{X_{i}}=(S_{i},\sigma_{i},A_{i},c_{i},\mathcal{D}_{i},V_{i},T_{i})$ for each variable $X_{i}$; and a feasibility constraint $\mathcal{F}\subseteq 2^{[n]}$. The constraint $\mathcal{F}$ corresponds to an upwards closed set in the minimization version (henceforth, min-CICS), and to a downwards closed set in the maximization version (henceforth, max-CICS). We use $\mathbb{M}:=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n})$ to denote the set of MDPs and $\mathbb{I}=(\mathbb{M},\mathcal{F})$ to denote the CICS instance.
General Policies for CICS.
@@
A policy (a.k.a. algorithm) for CICS proceeds as follows. Let $S\subseteq[n]$ denote the set of indices of all terminated MDPs. Let $s_{i}$ denote the current state of the MDP $\mathcal{M}_{i}$ at any point of time during the process. Initially, $S=\emptyset$ and $s_{i}=\sigma_{i}$ for all $i\in[n]$. The algorithm chooses at every step an index $i\in[n]\setminus S$ corresponding to a non-terminated MDP $\mathcal{M}_{i}$ and an action $a_{i}\in A_{i}(s_{i})$. It then follows the action at a cost of $c_{i}(a_{i})$. If $a_{i}$ is the accept action (i.e. $s_{i}\in T_{i}$ is a terminal state), it adds $i$ to $S$ and collects the value $v_{i}(s_{i})$. Otherwise, it updates the state of $\mathcal{M}_{i}$ to a new state drawn from the distribution $\mathcal{D}(s_{i},a_{i},\cdot)$, while the states of all other MDPs $\mathcal{M}_{i^{\prime}}$ with $i^{\prime}\neq i$ remain unchanged. Observe that the algorithm can make both of these choices – the index of the MDP to move in and the action to take in that MDP – adaptively, depending on the evolution of all $n$ MDPs in previous steps.
@@
We use $\mathrm{OPT}(\mathbb{I})$ to denote the cost/utility of the optimal policy for the CICS instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$.
Committing Policies for CICS.
A commitment $\pi_{i}$ for the MDP $\mathcal{M}_{i}$ is a mapping from every state $s_{i}\in S_{i}$ of $\mathcal{M}_{i}$ to a distribution $\pi_{i}(s_{i})$ over the actions in $A_{i}(s_{i})$. A committing policy is then characterized by a tuple of commitments $\mathcal{P}=(\pi_{1},\cdots,\pi_{n})$; at each step, the policy selects an MDP $\mathcal{M}_{i}$ at current state $s_{i}$ and takes an action that is sampled from $\pi_{i}(s_{i})$. Formally, for $i\in[n]$, let $\mathcal{M}_{i}^{\pi_{i}}$ denote the Markov chain resulting from applying the commitment $\pi_{i}$ to MDP $\mathcal{M}_{i}$. Then any policy for the instance $\mathbb{I}_{|\mathcal{P}}:=(\mathcal{M}_{1}^{\pi_{1}},\cdots,\mathcal{M}_{n}^{\pi_{n}},\mathcal{F})$ is a committing policy for $\mathbb{I}$ consistent with the tuple $\mathcal{P}$. Note that while committing policies must make all local decisions as dictated by $\mathcal{P}$, the index of the MDP to move in can be selected adaptively: committing policies are therefore adaptive algorithms.
We use $\mathcal{C}\left(\mathcal{M}_{i}\right)$ to denote the set of all possible commitments $\pi_{i}$ for the MDP $\mathcal{M}_{i}$ and $\mathcal{C}\left(\mathbb{I}\right)=\prod_{i}\mathcal{C}\left(\mathcal{M}_{i}\right)$ to denote the set of all possible tuples of commitments $\mathcal{P}$ for the CICS instance $\mathbb{I}$. The commitment gap quantifies the performance loss of committing policies relative to the optimal policy.
Definition 3 (Commitment Gap).
The commitment gap for any min-CICS (or max-CICS) instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ is defined as
$\mathrm{CG}(\mathbb{I}):=\min_{\mathcal{P}\in\mathcal{C}\left(\mathbb{I}\right)}\frac{\mathrm{OPT}(\mathbb{I}_{|\mathcal{P}})}{\mathrm{OPT}(\mathbb{I})}\geq 1\;\;\;\;\Big{(}\text{or }\;\mathrm{CG}(\mathbb{I}):=\max_{\mathcal{P}\in\mathcal{C}\left(\mathbb{I}\right)}\frac{\mathrm{OPT}(\mathbb{I}_{|\mathcal{P}})}{\mathrm{OPT}(\mathbb{I})}\leq 1\Big{)}.$
Matroid-min-CICS.
From now on, we focus primarily on matroid feasibility constraints: each MDP corresponds to an element of some known matroid $\mathrm{M}=([n],\mathcal{I})$. For min-CICS, $\mathcal{F}$ is the collection of all sets that contain a basis of $\mathrm{M}$; for max-CICS, $\mathcal{F}$ contains all the independent sets of $\mathrm{M}$. Observe that single-element selection, where the algorithm’s goal is to accept one MDP, is a special case of matroid selection. Following the work of Singla (2017), our results will apply to any feasibility constraint that admits an efficient frugal approximation algorithm. We describe this extension in Appendix B. For conciseness, in the main body of the paper, we state our results for the minimization setting (matroid-min-CICS); the analogous framework and results for the maximization setting are discussed in Appendix A.